X Company wants to estimate total costs (overhead plus administrative) in March, when production is expected to be 1,300 units. It analyzed January costs to determine the parameters of the overhead and administrative cost functions. January production was 1,000 units and costs were as follows:

Total overhead costs $59,000
Total fixed overhead costs 33,040
Total administrative costs 34,000
Total variable administrative costs 8,500


What is X Company's estimate of total costs in March [round unit costs to two decimal places]?
Answer #1

Total overhead costs = $59,000

Total fixed overhead costs = $33,040

Total variable overhead = Total overhead costs - Total fixed overhead costs

= 59,000-33,040

= $25,960

Variable overhead cost per unit = Total variable overhead/ Number of units

= 25,960/1,000

= $25.96

Total administrative costs = $34,000

Total variable administrative costs = $8,500

Total fixed administrative cost = Total administrative costs -Total variable administrative costs

= 34,000-8,500

= $25,500

Total variable administrative cost per unit = 8,500/1,000

= $8.5

Calculation of Total Costs in March
Total variable overhead cost (25.96 x 1,300) 33,748
Total fixed overhead cost 33,040
Total variable administrative costs (8.5 x 1,300) 11,050
Total fixed administrative cost 25,500
Total Cost $103,338

Kindly comment if you need further assistance. Thanks‼!